\begin{thebibliography}{12}
\bibliographystyle{alpha}
\bibitem{M. Aigner}  M. Aigner. Motzkin numbers. European Journal of Combinatorics, 19(6) (1998) 663--675.
\bibitem{M. Aigner Catalan-like} M. Aigner. Catalan-like numbers and determinants, J. Combin. Theory Ser. A 87 (1999) 33--51.
\bibitem{M. Aigner Catalan} M. Aigner. Catalan and other numbers: a recurrent theme, in: H. Crapo, D. Senato (Eds.), Algebraic Combinatorics and Computer Science, Springer, Berlin, 2001, pp. 347--390.

\bibitem{P. Barry} P. Barry. Some conjectures on the ratio of Hankel transforms for sequences and series reversion, 17 Jan. 2007. arxiv:math/0701483v1[math.CO].
\bibitem{P. Barry Catalan}  P. Barry. Generalized Catalan numbers, Hankel transforms and Somos-4 sequences, J. Integer Seq. 13 (2010) Article 10.7.2.

\bibitem{D.M. Bressoud}  D. M. Bressoud. Proofs and Confirmations: The story of the alternating sign matrix conjecture. Cambridge University Press Cambridge, England, 1999.

\bibitem{R. A Brualdi and S. Kirkland.}  R. A Brualdi and S. Kirkland. Aztec diamonds and digraphs, and Hankel determinants of schr\"{o}der numbers. J.  Combin. Theory, Series B, 94(2) (2005) 334--351.

\bibitem{X. K. Chang-X. B. Hu and G. Xin} X. K. Chang, X. B. Hu, and G. Xin. Hankel determinant solutions to several discrete integrable systems and the Laurent property, SIAM J. Disc. Math., 29(1) (2015) 667--682.

\bibitem{Chien-Eu-Fu}Hsu-Lin Chien, Sen-Peng Eu and Tung-Shan Fu, On Hankel determinants for Dyck paths with peaks avoiding multiple classes of heights, European Journal of Combinatorics, 101(2022), 103478.
    
\bibitem{J.Cigler-Catalan} J. Cigler. Catalan numbers, Hankel determinants and Fibonacci polynomials. Arxiv preprint arXiv:1801.05608, 2018.

\bibitem{J.Cigler} J. Cigler. Some nice Hankel determinants. Arxiv:1109.1449, 2011.

\bibitem{J.Cigler-conjecture} J. Cigler. Some remarks and conjectures about Hankel determinants of polynomials which are related to Motzkin paths. Arxiv:2204.09910, 2022.
    
\bibitem{Ciglerconvolution} J. Cigler, Hankel determinants of convolution powers of Catalan numbers revisited, arXiv:2403.11244.
    

%\bibitem{J.Cigler-C.Krattenthaler}  J. Cigler and C. Krattenthaler. Some determinants of path generating functions. Adv. Appl. Math., 46(1) (2011) 144--174.
%
%\bibitem{J.Cigler-Mike}   J. Cigler and M. Tyson. An interesting class of Hankel determinants.  Arxiv preprint arXiv:1807.08330, 2018.
%
%\bibitem{Omer Egecioglu} \"{O}mer E\u{g}ecio\u{g}lu, Timothy Redmond and Charles Ryavec. A multilinear operator
%for almost product evaluation of Hankel determinants. Journal of Combinatorial
%Theory, Series A, 117(1) (2010) 77--103.
%
%
%\bibitem{Omer Egecioglu 0} \"{O}mer E\u{g}ecio\u{g}lu, Timothy Redmond and Charles Ryavec.
%Almost product evaluation of Hankel determinants, Electron. J. Combin. 15 (1) (2008)
%\#R6, 58 pp.
\bibitem{M. Elouafi} M. Elouafi. A unified approach for the Hankel determinants of classical combinatorial numbers, J. Math. Anal. Appl. 431 (2015) 1253--1274.
    
\bibitem{I. Gessel and G. Viennot.}  I. Gessel and G. Viennot. Binomial determinants, paths, and hook length formulae. Adv. in Math., 58(3) (1985) 300--321.


\bibitem{Fulmek} M. Fulmek, Hankel Determinants of convoluted Catalan numbers and nonintersecting lattice paths: A bijective proof of Cigler's Conjecture, arXiv:2402.19127.
    
\bibitem{Gessel and Xin}  I. M. Gessel and G. Xin. The generating function of ternary trees and continued fractions. Electronic J. Combin. 13 (2008) \# R53.
    
\bibitem{Q.-H. Hou-A. Lascoux-Y.-P. Mu} Q.-H. Hou, A. Lascoux, Y.-P. Mu. Evaluation of some Hankel determinants, Adv. Appl. Math. 34 (2005) 845--852.

\bibitem{W. B. Jones and W. J. Thron}  W. B. Jones and W. J. Thron. Continued Fractions: Analytic Theory and Applications, Encyclopedia of Mathematics and its Applications. vol. 11, Addison-Wesley Publishing Co., Reading, Mass., 1980.


\bibitem{C. Krattenthaler.1999}  C. Krattenthaler. Advanced determinant calculus. S\'{e}minaire Lotharingien Combin., 42 (1999) B42q.

\bibitem{C. Krattenthaler.}   C. Krattenthaler. Advanced determinant calculus: a complement. Linear Algebra Appl., 411:68--166, 2005.

\bibitem{C. Krattenthaler.2010} C. Krattenthaler. Determinants of (generalised) Catalan numbers, J. Statist. Plann. Inference 140 (2010) 2260--2270.

%\bibitem{C.Krattenthaler-D.Yaqubi} C. Krattenthaler and D. Yaqubi. Some determinants of path generating functions, II. Arxiv preprint arXiv:1802.05990.

\bibitem{MTM2016}  C. Mariconda, A. Tonolo, C. Mariconda, and A. Tonolo. Linear Recurrence Relations. Discrete Calculus: Methods for Counting, 2016, 355-417.

\bibitem{J.M.E.Mays J.Wojciechowski}  M. E. Mays and J. Wojciechowski. A determinant property of Catalan numbers. Discrete Math., 211(1-3) (2000) 125--134.


\bibitem{Mu.Lili-Wang.Yi-Yeh.Yeong-Nan}L. Mu,  Y. Wang and Y. Y. Nan. Hankel determinants of linear combinations of consecutive Catalan-like numbers. Discrete Math. 340 (2017), no. 12, 3097--3103.

\bibitem{Mu.Lili-Wang.Yi}L. Mu and Y. Wang. Hankel determinants of shifted Catalan-like numbers. Discrete Math. 340 (2017), no. 6, 1389--1396.

\bibitem{R.Sulanke and Xin} R. Sulanke and G. Xin. Hankel determinants for some common lattice paths. Adv. Appl. Math. 40 (2008) 1410--167.

\bibitem{U. Tamm}U. Tamm. Some aspects of Hankel matrices in coding theory and combinatorics, Electron. J. Combin. 8 (1) (2001) 31.


\bibitem{Y. Wang and G. Xin}Y. Wang, G. Xin and M. Zhai. Hankel determinants and shifted periodic continued fractions, Adv. Appl. Math.
102 (2019) 83--112.


\bibitem{Wang-Zhang} Y.  Wang, Z. Zhang, Sufficient condition for $(\alpha, \beta)$ Somos 4Hankel determinants, Discrete Mathematics, 347(2024), 113937.

\bibitem{Xin Somos4}  G. Xin. Proof of the Somos-4 Hankel determinants conjecture. Adv. Appl. Math. 42 (2009) 152--156.








\end{thebibliography}

\documentclass[reqno, 12pt]{amsart}

\pdfoutput=1

\usepackage{latexsym}
\usepackage[centertags]{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{newlfont}
\usepackage{graphics}
\usepackage{color}
\usepackage{float}
\usepackage{diagbox}
\usepackage{longtable}
\usepackage{rotating}
\usepackage{multirow}

\usepackage{extarrows}

\usepackage[sort,compress,numbers]{natbib}

\usepackage[utf8]{inputenc}

\newtheorem{thm}{Theorem}
\newtheorem{cor}[thm]{Corollary}
\newtheorem{lem}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}

\theoremstyle{mydefinition}

\newtheorem{exa}[thm]{Example}

\theoremstyle{myremark}

\allowdisplaybreaks[1]

\def\pa[1]{\frac{\partial}{\partial x}}

\newcommand{\qfac}[1]{(q)_n}

\title{Hankel Determinants for a Class of Weighted Lattice Paths}

\author{Ying Wang$^1$ and Zihao Zhang$^{2,*}$}

\address{ $^{1}$School of Mathematics and Statistics, North China University of Water Resources and Electric Power,
 Zhengzhou 450045, PR China
 }

 \address{$^{2}$School of Mathematical Sciences, Capital Normal University,
 Beijing 100048, PR China
 }

\email{$^1$\texttt{wangying2019@ncwu.edu.cn}\ \ \ \& $^2$\texttt{zihao-zhang@foxmail.com}}
\date{\today}

\begin{document}

\maketitle

\begin{abstract}
 In this paper, our primary goal is to calculate the Hankel determinants for a class of lattice paths, which are distinguished by the step set consisting of \(\{(1,0), (2,0), (k-1,1), (-1,1)\}\), where the parameter \(k\geq 4\). These paths are constrained to return to the $x$-axis and remain above the \(x\)-axis.
 When calculating for \(k = 4\), the problem essentially reduces to determining the Hankel determinant of \(E(x)\), where \(E(x)\) is defined as
\[ E(x) = \frac{a}{E(x)x^2(dx^2 - bx - 1) + cx^2 + bx + 1}. \]
Our approach involves employing the Sulanke-Xin continued fraction transform to derive a set of recurrence relations, which in turn yield the desired results. For \(k \geq 5\), we utilize a class of shifted periodic continued fractions as defined by Wang-Xin-Zhai, thereby obtaining the results presented in this paper.

\end{abstract}

\noindent
\begin{small}
 \emph{Mathematic subject classification}: Primary 15A15; Secondary 05A15, 11B83.
\end{small}

\noindent
\begin{small}
\emph{Keywords}: Hankel determinants; Lattice path; Continued fractions.
\end{small}

\newcommand\con[1]{\equiv_{#1}}

\section{Introduction}

This paper delves into the application of generating functions to the computation of Hankel determinants. Specifically, we define the Hankel determinant of a generating function \( A(x) = \sum_{n \geq 0} a_n x^n \) as
\[ H_n(A(x)) = \det (a_{i+j})_{0 \leq i, j \leq n-1}, \]
with the convention that \( H_0(A(x)) = 1 \).

There has been a proliferation of research dedicated to the evaluation of Hankel determinants corresponding to a variety of lattice path counting numbers. Classical lattice path Hankel determinants include binomial coefficients, Catalan numbers \cite{J.M.E.Mays J.Wojciechowski} and their shifts, Motzkin numbers \cite{M. Aigner,J.Cigler,J.Cigler-conjecture}, large and little Schröder numbers \cite{R. A Brualdi and S. Kirkland.}, among others. Special cases such as peak-height-restricted lattice paths \cite{Chien-Eu-Fu,Ciglerconvolution}, weighted Dyck paths \cite{M. Aigner Catalan-like}, and those without closed forms \cite{X. K. Chang-X. B. Hu and G. Xin,P. Barry, Xin Somos4, Wang-Zhang} have also been studied. For further references, see \cite{M. Aigner Catalan-like, M. Aigner Catalan, P. Barry,P. Barry Catalan, D.M. Bressoud, R. A Brualdi and S. Kirkland., M. Elouafi, I. Gessel and G. Viennot., Q.-H. Hou-A. Lascoux-Y.-P. Mu,C. Krattenthaler.1999, C. Krattenthaler.,C. Krattenthaler.2010,Mu.Lili-Wang.Yi-Yeh.Yeong-Nan,Mu.Lili-Wang.Yi, R.Sulanke and Xin, U. Tamm}.

The classical method of continued fractions, either by \(J\)-fractions (Krattenthaler \cite{C. Krattenthaler.}) or by \(S\)-fractions (Jones and Thron \cite[Theorem 7.2]{W. B. Jones and W. J. Thron}), requires \(H_n(A(x)) \neq 0\) for all \(n\). Gessel-Xin's \cite{Gessel and Xin} continued fraction method, however, allows \(H_n(A(x)) = 0\) for some values of \(n\). Their method is based on three rules about two-variable generating functions that can transform one set of determinants to another set of determinants of the same values. These rules correspond to a sequence of elementary row or column operations. This method was systematically used by Sulanke-Xin \cite{R.Sulanke and Xin} for evaluating Hankel determinants of quadratic generating functions, such as known results for Catalan numbers, Motzkin numbers, Schröder numbers, etc.

Sulanke-Xin defined a quadratic transformation \(\tau\) such that \(H(F(x))\) and \(H(\tau(F(x)))\) have simple connections. Recently, shifted periodic continued fractions (of order \(q\)) of the form
$$
F_0^{(p)}(x) \mathop{\longrightarrow}\limits^\tau F_1^{(p)}(x)\mathop{\longrightarrow}\limits^\tau \cdots \mathop{\longrightarrow}\limits^\tau F_q^{(p)}(x)=F_0^{(p+1)}(x)
$$
were found in \cite{Y. Wang and G. Xin}, which appear in Hankel determinants of many path counting numbers. Here \(p\) is an additional parameter. If one can guess an explicit formula for \(F_0^{(p)}(x)\), then their Hankel determinants can be easily computed.

For the lattice paths defined by the path set \(\{(1,0),(k-1,1),(-1,1)\}\),  where \(k \geq 2\).
The generating function for these cases is expressed as follows:
$$
G^{1,k}(x,a,b_1,c) = \frac{a}{1 + b_1 x + cx^k G^{1,k}(x,a,b_1,c)}.
$$
The case where \( k = 2 \) corresponds to the weighted Motzkin numbers and its Hankel determinants has been investigated. For \( k > 2 \), the Hankel determinants are provided in \cite{J.Cigler, Y. Wang and G. Xin},

In this paper, our primary aim is to compute the Hankel determinants of lattice paths defined by the path set \(\{(1,0),(2,0),(k-1,1),(-1,1)\}\), where \(k \geq 4\). The weights associated with these paths are \(b_1\) for \((1,0)\),    \(b_2 \neq 0\) for \((2,0)\), \(c\) for \((k-1,1)\), and \(1\) for \((-1,1)\). The generating function is given by
$$
G^{2,k}(x,a,b_1,b_2,c) = \frac{a}{1 + b_1x + b_2x^2 + cx^k G^{2,k}(x,a,b_1,b_2,c)}.
$$

For \(k=1,2,3\), the formulas of  $H_n(G^{2,k}(x,a,b_1,b_2,c))$  are typically available only for a few certain special weights. For instance, when \(k=3\), the formula of $H_n(G^{2,k}(x,a,t,t+1,1))$ can be found in \cite{Y. Wang and G. Xin}.


In general, for \(r>2\) and \(k > 2r\), the path set \(\{(1,0),(2,0),\dots,(r,0), (k-1,1),(-1,1)\}\), the calculation process is analogous to that presented in this paper.
A specific instance is resolvable for particular values of $k = 2r + 1$ or $k = 2r$, exemplified by the convolution of Catalan numbers \cite{J.Cigler-Catalan,Ciglerconvolution,Fulmek}. However, it is challenging to compute when $k \leq 2r$ even for a fixed $r$.

In this paper, we also focus on the continued fraction $E(x)$ which satisfied
\begin{equation}
E(x) = \frac{a}{E(x)x^2(dx^2 - bx - 1) + cx^2 + bx + 1}, \label{eq:E_x}
\end{equation}
which arise from the process of calculating $G^{2,4}(x,a,b_1,b_2,c)$.  Furthermore, when $a=b=d=1$, $c=-1$, the Hankel determinants of  $E(x)$ is the odd Fibonacci repeated (OEIS.A094967). See the proof
in the Section \ref{sec:recurrences}.

For convenient, we denote $B_n(a,c,d)$ as a second-order  linear homogeneous recurrence relation defined by
$$ B_{n+1} (a,c,d) = \beta B_n (a,c,d) + \alpha B_{n-1} (a,c,d) $$
with initial conditions $B_0 =1 $ and $B_1 = c - a$, and constants $\alpha = a(c + d - a)$, $\beta = c - 2a$.
It is well-established that the second-order  linear homogeneous recurrence have been extensively investigated. For a relevant reference, see \cite{MTM2016}.

The main results of this paper are as follows:

\begin{thm}\label{thm-e}
If $H_{n}(E)\neq0$ for all $n$, then we have
\begin{align*}
  H_{2n+1}(E) &=a^{(n+1)^2} (c + d - a)^{n^2}B_n(a,c,d),\\
  H_{2n}(E)  &=(-1)^na^{n(n+1)}  (c + d - a) ^{n(n-1)}B_n(a,c,d),
\end{align*}
\end{thm}
For \(k \geq 4\), we have the following two theorem about the Hankel determinants of \(G^{2,k}(x,a,b_1,b_2,c)\)
\begin{thm}\label{th-g24}
  If $H_n(G^{2,4}(x,a,b_1,b_2,c)) \neq 0$ for all $n\geq 0$. Then
  \begin{align*}
    H_{2n+1}(G^{2,4}(x,a,b_1,b_2,c)) =& b_2 a^{2n+1} (ac)^{n^2} B_{n-1}\left(\frac{ac}{b_2}, \frac{2ac+b_2^2}{b_2}, -\frac{ac}{b_2}\right),\\
     H_{2n}(G^{2,4}(x,a,b_1,b_2,c))=&(-1)^nb_2a^{2n}(ac)^{n(n-1)} B_{n-1}\left(\frac{ac}{b_2}, \frac{2ac+b_2^2}{b_2}, -\frac{ac}{b_2}\right).
  \end{align*}
\end{thm}

\begin{thm}\label{thmG2k}
When $k \geq 5$, we have
\begin{align*}
 H_{ k  n }(G^{2,k}(x,a,b_1,b_2,c)) &=\mu_1{a}^{ k  \,n \left( n-1 \right)+(2k-1)\,n}{c}^{ k  \,n \left( n-1 \right) +(k-1)\,n},\\
  H_{ k  n+1}(G^{2,k}(x,a,b_1,b_2,c)) &=\lambda_1{a}^{ k  \,n \left( n+1 \right)+n+1 }{c}^{ k  \,n
 \left( n+1 \right) -(k-1)\,n} , \\
  H_{ k  n+2}(G^{2,k}(x,a,b_1,b_2,c)) &=\lambda_2 \left( n+1 \right) b_{{2}}{a}^{ k  \,n \left( n+1 \right)+3\,n+2 }{c}^{ k  \,n \left( n+1 \right) -(k-3)\,n}, \\
  H_{ k  n+3}(G^{2,k}(x,a,b_1,b_2,c)) &=\cdots= H_{ k  n+k-2}(G^{2,k}(x,a,b_1,b_2,c)) =0,\\
 H_{ k  n+k-1}(G^{2,k}(x,a,b_1,b_2,c)) &=\mu_2\left( n+1 \right) b_{{2}}{a}^{ k  \,n \left( n+1 \right)+(2k-3)\,n+(2k-4) }{c}^{ k  \,n \left( n+1 \right) +(k-3)\,n+(k-3)}.
\end{align*}
For $k $ mod 4=0 or 1: we have $\lambda_1=-\lambda_2=\mu_1=\mu_2=1.$\\
For $k $ mod 4=2 : we have $\lambda_1=-\lambda_2=(-1)^n, \mu_1=\mu_2=\left( -1 \right) ^{n+k+1}.$\\
For $k $ mod 4=3 : we have $\lambda_1=-\lambda_2=(-1)^n, \mu_1=\mu_2=\left( -1 \right) ^{n+k}.$
\end{thm}


The paper is organized as follows. Section \ref{maintool} present the primary method for solving
Hankel determinants, namely the continued fraction approach of Sulanke and Xin.
Section \ref{sec:recurrences} provide several recurrence relations that are utilized in the computation
of $G^{2,4}(x,a,b_1,b_2,c)$.
In Section \ref{systemg24} , we evaluate the Hankel determinant of $G^{2,4}(x,a,b_1,b_2,c)$.
Section \ref{systemg2r} proof the Hankel determinant for \(G^{2,k}(x,a,b_1,b_2,c)\) for $k>4$.

\section{Main tool}\label{maintool}

We will introduce the continued fraction method of Sulanke and Xin, especially their quadratic transformation $\tau$ in \cite{R.Sulanke and Xin}.
This is the main tool of this paper.

\subsection{Sulanke-Xin's quadratic transformation $\tau$\label{sec:sulanke-xin-trans}}
This subsection is copied from \cite{Y. Wang and G. Xin}. We include it here for the reader's convenience.

Suppose the generating function $F(x)$ is the unique solution of a quadratic functional equation which can be written as
\begin{gather}
  F(x)=\frac{x^d}{u(x)+x^kv(x)F(x)},\label{xinF(x)}
\end{gather}
where $u(x)$ and $v(x)$ are rational power series with nonzero constants, $d$ is a nonnegative integer, and $k$ is a positive integer.
We need the unique decomposition of $u(x)$ with respect to $d$: $u(x)=u_L(x)+x^{d+2}u_H(x)$ where $u_L(x)$ is a polynomial of degree at most $d+1$ and $u_H(x)$ is a power series.
Then Propositions 4.1 and 4.2 of \cite{R.Sulanke and Xin} can be summarized as follows.
\begin{prop}\label{xinu(0,i,ii)}
Let $F(x)$ be determined by  \eqref{xinF(x)}. Then the quadratic transformation $\tau(F)$ of $F$ defined as follows gives close connections
between   $H(F)$ and $H(\tau(F))$.
\begin{enumerate}
\item[i)] If $u(0)\neq1$, then $\tau(F)=G=u(0)F$ is determined by $G(x)=\frac{x^d}{u(0)^{-1}u(x)+x^ku(0)^{-2}v(x)G(x)}$, and $H_n(\tau(F))=u(0)^{n}H_n(F(x))$;

\item[ii)] If $u(0)=1$ and $k=1$, then $\tau(F)=x^{-1}(G(x)-G(0))$, where $G(x)$ is determined by
$$G(x)=\frac{-v(x)-xu_L(x)u_H(x)}{u_L(x)-x^{d+2}u_H(x)-x^{d+1}G(x)},$$
and we have
$$H_{n-d-1}(\tau(F))=(-1)^{\binom{d+1}{2}}H_n(F(x));$$

\item[iii)] If $u(0)=1$ and $k\geq2$, then $\tau(F)=G$, where $G(x)$ is determined by
$$G(x)=\frac{-x^{k-2}v(x)-u_L(x)u_H(x)}{u_L(x)-x^{d+2}u_H(x)-x^{d+2}G(x)},$$
and we have
$$H_{n-d-1}(\tau(F))=(-1)^{\binom{d+1}{2}}H_n(F(x)).$$\label{xinu(0)(ii)}

\end{enumerate}

\end{prop}

\section{The Solution of a System of Recurrences} \label{sec:recurrences}
In this chapter, we compute the Hankel determinants of the continued fractions $E(x)$:
\begin{equation}
E(x) = \frac{a}{E(x)x^2(dx^2 - bx - 1) + cx^2 + bx + 1}, \label{eq:E_x}
\end{equation}
which will  be shown to emerge naturally in the next section when calculating the Hankel determinants related to $G^{2,4}(x,a,b_1,b_2,c)$.

Indeed, when applying the Proposition \ref{xinu(0,i,ii)} to $E(x)$, it is imperative to consider whether $a$ is zero. For expediency, we introduce the notation $\hat{a}$ to denote $a-c$.
We first analyze the case for $\hat{a} = a - c \neq 0$ and present the following lemma.
\begin{lem}\label{lem:e-f-g}
  Let $E(x)$ be as specified in  \eqref{eq:E_x} , and assume that $\hat{a} = a - c \neq 0$. Consider the continued functions $F(x)=\tau (E(x))$ and $G(x)=\tau^2 (E(x))$ :
  \begin{align}
    F(x) &= \frac{ad x^2 - b(a - c)x + a - c}{1 + bx - cx^2 - F(x)x^2}, \label{eq:F_x}\\
    G(x) &= \frac{\frac{a(a - c - d)}{a - c}}{G(x)x^2\left(\frac{ad}{a - c}x^2 - bx - 1\right) + (c - \frac{2ad}{a - c})x^2 + bx + 1}. \label{eq:G_x}
  \end{align}
Then  for all integers $n \geq 1$, the following relations hold:
    \[ H_n(E) = a^n H_{n-1}(F), \]
    and for $n \geq 2$,
    \[ H_n(E)= a^n (a - c)^{n-1} H_{n-2}(G) = a^n \hat{a}^{n-1} H_{n-2}(G). \]
\end{lem}

Our proof is by iterative application the above Lemma \ref{lem:e-f-g}. To be
precise, define $E_0(x)=E(x)$, and recursively define $E_{n}(x)$
to be the unique power series solution of
\begin{align}
E_{n+1}(x)={\frac {a_{n+1} }{E_{n+1}(x)~{x}^{2} \left( d_{n+1} {x}^{2
}-b_{n+1} x-1 \right) +c _{n+1} {x}^{2}+b_{n+1} x+1}}
\end{align}
where

\begin{align}
a_{n+1}&=\frac{a_{n}^2 -a_n c_n - a_n d_n}{a_n-c_n}\label{e-aa}\\
b_{n+1}&=b_{n}\label{e-bb}\\
c_{n+1}&={\frac {a_nc_n-2\,a_nd_n-{c_n}^{2}}{a_n-c_n}}\label{e-cc}\\
d_{n+1}&={\frac {a_nd_n}{a_n-c_n}}\label{e-dd}.
\end{align}

To rigorously  apply Lemma \ref{lem:e-f-g} for the evaluation of $H_n(E(x))$, it is imperative that $\hat{a}_k := a_k - c_k \neq 0$ and $a_k \neq 0$ for all $k \leq 2n+1$. Under this assumption, it can be readily verified that:
\begin{align}
     H_{2n+1}(E)=& a_0^{2n+1}\hat{a}_0^{2n} a_1^{2n-1} \hat{a}_1^{2n-2} \cdots a_{n-1}^3\hat{a}_{n-1}^2 a_{n}, \label{he2n+1} \\
      H_{2n}(E) =& a_0^{2n}\hat{a}_0^{2n-1} a_1^{2n-2} \hat{a}_1^{2n-3} \cdots a_{n-1}^2 \hat{a}_{n-1}. \label{he2n}
\end{align}
The fact that the recursion system can be resolved for any arbitrary initial condition comes as a surprise. Following this, we introduce the recurrence relations for \( a_n \) and \( \hat{a}_n \).

\begin{thm} \label{thm-rec-a}
   Suppose  $\hat{a}_k = a_k - c_k \neq0$  and $a_k \neq 0$  for $k\le n$, and the sequences $a_n$, $c_n$, and $d_n$ satisfy the recursions \eqref{e-aa}, \eqref{e-bb}, \eqref{e-cc}, and \eqref{e-dd}. Then the following relations hold for $n> 0$:
   \begin{align}
     a_{n+1}(c_0 - 2a_0 + a_n) &= a_0(c_0+d_0 -a_0), \label{eq:first_relation}\\
     \hat{a}_{n+1} & =- a_{n+1}+ 2a_0 - c_0 . \label{eahat}
   \end{align}
\end{thm}

\begin{proof}
   We aim to express all quantities in terms of the sequence $a_n$.
    Utilizing \eqref{e-cc} and the condition $a_n \neq c_n$, we have
    $$d_{n} = \frac{a_{n} c_{n}-a_{n} c_{n +1}-c_{n}^{2}+c_{n} c_{n +1}}{2 a_{n}}.$$
    Substituted $d_n$ to equations \eqref{e-aa} and \eqref{e-dd} following two equation respectively,
    \begin{align}
    2a_{n+1}- c_{n+1} -  ( 2a_n -c_n ) &= 0, \label{eq:recursed_first}\\
     c_{n+1}^2 - 2c_{n+1}a_{n+1} - c_{n+2}c_{n+1} + a_{n+1}c_n + c_{n+2}a_{n+1} &= 0. \label{e-r1}
   \end{align}
Under the condition $a_k \neq c_k$ for $k\le n$, the  equation \eqref{eq:recursed_first} is to say $$2a_{n+1}- c_{n+1} = 2a_0- c_0.$$
Using the notation $\hat a_{n+1} =a_{n+1} -c_{n+1}$, we obtain \eqref{eahat} and  $ c_{n+1} =2a_{n+1}-2a_0+ c_0$.
By substituting $ c_{n+1}$ to \eqref{e-r1}, we obtain
   \[(-c_0 + 2a_0 - a_{n+1})a_{n+2} + (c_0 - 2a_0 + a_n)a_{n+1} = 0.\]
   Under the condition $a_k \neq c_k$ for $k\le n$,
   this equation can also be recursively reduced to the initial values, leading to
   \[a_{n+1}(c_0 - 2a_0 + a_n) = c_0a_1 - a_0a_1.\]
Moveover by Lemma  \ref{lem:e-f-g}, we have $a_1 =\frac{a_0(a_0-c_0-d_0)}{a_0-c_0}$  and then
   we obtained \eqref{eq:first_relation}
   this completes the proof.
\end{proof}

Indeed, the recursion \eqref{eq:first_relation} and \eqref{eahat} can be converted to a solvable second-order linear homogeneous recurrence relation for its numerators or denominators.

\begin{cor}\label{cor-an}
Let $a_n = \frac{A_n}{B_n}$ for $n\geq 0$. It follows that
\[ a_{n+1} = \frac{\alpha B_{n}}{B_{n+1}} \quad \text{and} \quad \hat{a}_{n+1} = -\frac{B_{n+2}}{B_{n+1}}, \]
where $\alpha = a_0(c_0 + d_0 - a_0)$. Let $\beta = c_0 - 2a_0$. Then, the sequence $\{B_n\}$ satisfies the linear homogeneous recurrence relation
\[ B_{n+1} = \beta B_n + \alpha B_{n-1} \]
with initial values $B_0 = 1$ and $B_1 = c_0 - a_0$.
\end{cor}
\begin{proof}
We rewrite the equation \eqref{eq:first_relation}:
\[
a_{n+1} = \frac{\alpha}{\beta  + a_{n}}.
\]
Substituting the definition of $a_n$, we obtain
\[
\frac{A_{n+1}}{B_{n+1}} = \frac{\alpha}{\beta + \frac{A_{n}}{B_{n}}} = \frac{\alpha B_{n}}{\beta B_n + A_n}.
\]
This simplifies to the recurrence relations
\begin{align*}
  A_{n+1} &  =\alpha B_{n}, \\
 B_{n+1} &= \beta B_n + A_n =\beta B_{n} + \alpha B_{n-1}.
\end{align*}
It is easy to verify the initial values.

By \eqref{eahat}, it is easy to derive that
$\hat{a}_{n+1} =-\frac{A_{n+1}}{B_{n+1}}-\beta=\frac{-\beta B_n-\beta B_{n+1}}{B_{n+1}}=-\frac{B_{n+2}}{B_{n+1}} .$
\end{proof}

We are now prepared to demonstrate Theorem 1.
\begin{proof}[Proof of Theorem 1]
Assume that $H_{n} \neq 0$. According to Equations \eqref{he2n+1} and \eqref{he2n}, this condition is equivalent to stating that $a_n = 0$ and $\hat{a}_{n-1} = 0$, while $B_n \neq 0$.
In light of Corollary \ref{cor-an}, and utilizing Equations \eqref{he2n+1} and \eqref{he2n}, we proceed to make the necessary substitutions, which will allow us to establish Theorem 1.
\end{proof}

Here, we present a well-known general term formula for a second-order homogeneous linear recurrence relation.

\begin{prop}\label{pror-f}
The recurrence relation for the sequence $f_n$ is given by $f_n = \beta f_{n-1} + \alpha f_{n-2}$. Consequently,
if $\mu, \gamma$ are roots of the characteristic equation $x^2 - \beta x - \alpha = 0$,
$$
f_{n-1} =
\begin{cases}
\frac{\mu^{n-1} (f_1 - f_0 \gamma) - \gamma^{n-1}(f_1 - f_0 \mu)}{\mu - \gamma}, & \text{if } \mu \neq \gamma, \\
\mu^{n-2} ((n-1)f_1 - \mu (n-2) f_0), & \text{if } \mu = \gamma.
\end{cases}
$$
\end{prop}

%

\begin{exa}
We evaluate the Hankel determinants for the function $E(x)$, which satisfies the equation $E(x)= \frac{1}{E(x) \,x^{2} \left(x^{2}-x -1\right)-x^{2}+x +1}$.
By applying Theorem \ref{thm-e} and Proposition \ref{pror-f}, we obtain the following expressions for the Hankel determinants:
  \begin{eqnarray*}
   H_{2n}(E(x)) = H_{2n+1}(E(x)) = \frac{\left(-1\right)^{n^{2}}}{10} \left(\left(5+\sqrt{5}\right) \left(-\frac{3}{2}-\frac{\sqrt{5}}{2}\right)^{n}+\left(5-\sqrt{5}\right) \left(\frac{\sqrt{5}}{2}-\frac{3}{2}\right)^{n}\right).   \\
  \end{eqnarray*}
  It is a well-known fact that the $n-$th Fibonacci number $Fib(n)$ is equal to $$- \frac{1}{\sqrt{5}} \left(-\frac{\sqrt{5}}{2}+\frac{1}{2}\right)^{n}+\frac{1}{\sqrt{5}}  \left(\frac{\sqrt{5}}{2}+\frac{1}{2}\right)^{n}.$$
Through direct computation, we can confirm that $H_{2n}(E(x))= H_{2n+1}(E(x)) = Fib(2n+1)$.
\end{exa}


Upon repeated application of Lemma \ref{lem:e-f-g} to compute $H_n(E)$, if the condition arise where $a_k = c_k$ for the first $k < n$ within $E_k$, Lemma 5 becomes inapplicable. However, fortuitously, applying the $\tau$ transform four times leads to a return to the original form of $E(x)$. This is encapsulated in the following lemma:

\begin{lem}\label{lem:e-eac}
Assume $\hat{a} = a - c = 0$. The generating functions $E^i(x)$ and $F^j(x)$ for $i=0,1$ and $j=1,2,3$ are uniquely defined by the system:
\begin{align*}
    E^0(x) &= \frac{a}{E^0(x)x^2(dx^2 - bx - 1) + ax^2 + bx + 1},\\
    F^1(x) &= \frac{adx^2}{x^2F^1(x) + ax^2 - bx - 1},\\
    F^2(x) &= \frac{ad}{x^4F^2(x) + ax^2 - bx - 1},\\
    F^3(x) &= \frac{ad(x^2 - bx) - a}{F^3(x)x^2 - ax^2 - bx - 1},\\
    E^1(x) &= -\frac{d}{(x^2d - xb - 1)x^2E^1(x) - (a + 2d)x^2 + bx + 1}.
\end{align*}
Here, $F^j(x) = \tau^j(E^0(x))$ for $j=1,2,3$, and $E^1(x) = \tau^4(E^0(x))$. Then the relations hold:
\begin{align*}
    H_n(E^0) &= a^nH_{n-1}(F^1),\quad n \geq 1\\
    H_n(E^0) &= a^n(-ad)^{n-1}H_{n-4}(F^2), \quad n \geq 4,\\
    H_n(E^0) &= -a^n(ad)^{2n-5}H_{n-5}(F^3), \quad n \geq 5,\\
    H_n(E^0) &= -a^{2n-5}(ad)^{2n-5}H_{n-6}(E^1), \quad n \geq 6.
\end{align*}
\end{lem}

This Lemma suggests that such instances remain within the purview of our framework. Nevertheless, for arbitrary initial values $a_0, b_0, c_0, d_1$, determining the occurrence of $a_k = c_k$ necessitates a case-by-case analytical approach. See some examples in next section.

\section{ The Hankel Determinants of $G^{2,4}(x,a,b_1,b_2,c)$  }\label{systemg24}
\def\mT{\mathcal{T}}
For $r=2$, $k=4$,~$b_2 \neq 0 $,  we have the functional equation
\begin{align*}
  G^{2,4}(x,a,b_1,b_2,c)=\frac{a} { G^{2,4}(x,a,b_1,b_2,c)x^4 c+b_2 x^2 +b_1 x+1}.
\end{align*}

\begin{proof}[Proof of the Theorem \ref{th-g24}]
  We apply Proposition \ref{xinu(0,i,ii)} to $G_0:= G^{2,4}(x,a,b_1,b_2,c)$ by repeatedly using the transformation $\tau$.Then
\begin{align*}
  G_0(x)& \mathop{\longrightarrow}\limits^\tau G_1(x)\mathop{\longrightarrow}\limits^\tau G_{2}(x)
\end{align*}
where
 \begin{align*}
    G_1=\frac{-acx^2-b_2b_1 x-b_2}{-G_1 x^2-b_2 x^2+b_1 x+1}.
 \end{align*}
\begin{align*}
    G_2=\frac{\frac{ac}{b_2}}{G_2 x^2( -\frac{acx^2}{b_2}-b_1x-1)+ \frac{(2ac+b_2^2)x^2}{b_2}+b_1x+1}.
 \end{align*}
and
\begin{align}\label{24G0-G2}
 H_n(G_0) =a^n (-b_2)^{n-1}  H_{n-2}(G_2) .
\end{align}

By the assumption $H_n(G_2) \neq 0$ for all $n$, $G_2$ and $E$ share the same form.
By Theorem \ref{thm-e}, we have
\begin{align}
   H_{2n+1}(G_2) &=(ac)^{n^2}\left(\frac{ac}{b_2}\right)^{2n+1}B_n\left(\frac{ac}{b_2}, \frac{2ac+b_2^2}{b_2}, -\frac{ac}{b_2}\right), \label{G2n1} \\
    H_{2n}(G_2) &=(-1)^n(ac)^{n(n-1)}\left(\frac{ac}{b_2}\right)^{2n}B_n\left(\frac{ac}{b_2}, \frac{2ac+b_2^2}{b_2}, -\frac{ac}{b_2}\right) . \label{G2n}
\end{align}
Then the Theorem \ref{th-g24} follows by \eqref{24G0-G2},  \eqref{G2n1} and \eqref{G2n}.
\end{proof}

Firstly, Example is $H_n(G^{2,4}(x,a,b_1,b_2,c))\neq0$ for all $n$.
\begin{exa}
  Let $4ac+b_2^2=0$,
 this constitutes a system that we can solve by repeatedly using Lemma \ref{lem:e-f-g}.
 \begin{proof} Recalled that:
  $$ G^{2,4}(x,a,b_1,b_2,-\frac{b_2^2}{4a})={\frac {-4\,{a}^{2}}{G^{2,4}(x,a,b_1,b_2,-\frac{b_2^2}{4a})){x}^{4}{b_{{2}}}^{2}-4\,b_{{2}}{x}^{2}a-4\,b_{{1
}}xa-4\,a}}.$$
Then $G_2$ becomes
$$G_2=-{\frac {b_{{2}}}{ \left( b_{{2}}{x}^{4}-4\,b_{{1}}{x}^{3}-4\,{x}^{2}
 \right) G_2+2\,b_{{2}}{x}^{2}+4\,b_{{1}}x+4}}
$$
By Proposition \ref{pror-f}, we  have
$$B_{n-1}\left(-\frac{b_2}{4},\frac{b_2}{2},\frac{b_2}{4}\right) =\frac12\, \left( \frac12\,b_{{2}} \right) ^{n}+\frac14\,
 \left( 2\,n+2 \right)  \left( \frac12\,b_{{2}} \right) ^{n} \neq 0, \text{\ for all\ } n.
$$
By Theorem \ref{th-g24}, then we can obtain
\begin{align*}
  H_{2n+1}(G^{2,4}(x,a,b_1,b_2,-\frac{b_2^2}{4a})) =& b_2a^{2n+1}\left(-\frac{b_2^2}{4}\right)^{n^2}B_{n-1}\left(-\frac{b_2}{4},\frac{b_2}{2},\frac{b_2}{4}\right), \\
  H_{2n}(G^{2,4}(x,a,b_1,b_2,-\frac{b_2^2}{4a})) =&(-1)^n b_2a^{2n}\left(-\frac{b_2^2}{4}\right)^{n(n-1)}B_{n-1}\left(-\frac{b_2}{4},\frac{b_2}{2},\frac{b_2}{4}\right).
\end{align*}
 \end{proof}
\end{exa}
When $H_n(G^{2,4}(x,a,b_1,b_2,c))=0$ for a particular $n$, we have not identified a uniform formula for the Hankel determinants, and their periods remain indeterminable. To clarify this point, we offer three straightforward examples where $k ac + b^2 = 0$ for $k = 1, 2, 3$.

Example \ref{exa-ac} relies solely on Lemma \ref{lem:e-eac}, and the determinant exhibits a periodicity of six.
 \begin{exa}
 \label{exa-ac}
   When $ac+b_2^2=0$, we have
 \begin{align*}
   H_{6n}(G^{2,4}(x,a,b_1,b_2,-\frac{b_2^2}{a})) &= a^{6n}b_2^{18n^2-3n}, \\
   H_{6n+1}(G^{2,4}(x,a,b_1,b_2,-\frac{b_2^2}{a})) &= a^{6n+1}b_2^{18n^2+3n}, \\
   H_{6n+2}(G^{2,4}(x,a,b_1,b_2,-\frac{b_2^2}{a})) &=-a^{6n+2}b_2^{18n^2+9n+1}, \\
   H_{6n+3}(G^{2,4}(x,a,b_1,b_2,-\frac{b_2^2}{a})) &=-a^{6n+3}b_2^{18n^2+15n+3}, \\
   H_{6n+4}(G^{2,4}(x,a,b_1,b_2,-\frac{b_2^2}{a})) &=H_{6n+5}(G^{2,4}(x,a,b_1,b_2,-\frac{b_2^2}{a}))=0.
 \end{align*}
 \end{exa}
 \begin{proof}
   For  $ac+b_2^2=0$, then $G_2=-{\frac {b_{{2}}}{G_2{x}^{2} \left( b_{{2}}{x}^{2}-b_{{1}}x-1 \right) -b_{{2}}{x}^{2}+b_{{1}}x +1 } }$.
By using the Lemma \ref{lem:e-eac}.   This results in a periodic continued fractions of order 4 :
\begin{align*}
  G_2(x)& \mathop{\longrightarrow}\limits^\tau G_3(x)\mathop{\longrightarrow}\limits^\tau G_4(x)\mathop{\longrightarrow}\limits^\tau G_{5}(x)\mathop{\longrightarrow}\limits^\tau G_6(x)= G_2(x)\cdots.
\end{align*}
We obtain
\begin{align*}
  H_k(G_2)&=(-b_2)^kH_{k-1}(G_3), \quad  H_{k-1}(G_3)=b_2^{2(k-1)}H_{k-4}(G_4)\\
  H_{k-4}(G_4)&=b_2^{2(k-4)}H_{k-5}(G_5),\quad  H_{k-5}(G_5)=(-b_2)^{k-5}H_{k-6}(G_2).
\end{align*}
Combine them, we can get
$$ H_k(G_2)=-b_2^{6k-15}H_{k-6}(G_2).$$

The initial values are
\begin{align*}
  H_0(G_2)&=1,\ \ H_1(G_2)=(-b_2), \quad  H_2(G_2)=0,\ \  H_3(G_2)=0, \\
   H_{4}(G_2)&=b_2^{10}, \ \   H_{5}(G_2)=-b_2^{15}.
\end{align*}
Then we can get the result.
 \end{proof}

Example \ref{exa-2ac} relies on Lemma \ref{lem:e-f-g} and \ref{lem:e-eac}, and the determinant exhibits a periodicity of eight.
\begin{exa}
\label{exa-2ac}
  Let $2ac+b_2^2=0$, then $G_2$ becomes
$$G_2=-{\frac {b_{{2}}}{ G_2x^2\left( b_{{2}}{x}^{2}-2\,b_{{1}}{x}-2 \right)+2\,b_{{1}}x+2}}.
$$
The result is a periodic continued fractions of order 6.
\begin{align*}
   \underbrace{ G_2(x) \mathop{\longrightarrow}\limits^\tau G_3(x)\mathop{\longrightarrow}\limits^\tau G_4 }_{\text{Lemma \ref{lem:e-f-g}}}\underbrace{(x)\mathop{\longrightarrow}\limits^\tau G_{5}(x)\mathop{\longrightarrow}\limits^\tau G_6(x)\mathop{\longrightarrow}\limits^\tau G_7(x)\mathop{\longrightarrow}\limits^\tau G_8(x)}_{\text{Lemma \ref{lem:e-eac}}}= G_2(x)\cdots.
\end{align*}

We obtain
\begin{align*}
  H_k(G_2)&=\left(-\frac12b_2\right)^kH_{k-1}(G_3),\, \, H_{k-1}(G_3)=\left(-\frac12b_2\right)^{k-1}H_{k-2}(G_4), \\
  H_{k-2}(G_4)&=\left(-b_2\right)^{k-2}H_{k-3}(G_5),\, \,H_{k-3}(G_5)=\left(\frac12b_2\right)^{2k-6}H_{k-6}(G_6),\\
  H_{k-6}(G_6)&=\left(\frac12b_2\right)^{2k-12}H_{k-7}(G_7),\, \, H_{k-7}(G_7)=\left(-b_2\right)^{k-7}H_{k-8}(G_2).
\end{align*}
Combine them, then we can get
$$H_{k}(G_2)=\left(\frac12\right)^{6k-19}b_2^{8k-28} H_{k-8}(G_2).$$
By using some initial values,  one can readily obtain the Hankel determinants (we omit here).
\end{exa}

\begin{exa}
  Consider the case where $3ac + b_2^2 = 0$. We examine a periodic continued fraction of order 10, and observe that the period of the sequence $H_{n}(G^{2,4}(x,a,b_1,b_2,-\frac{b_2^2}{3a}))$ is 12. Within each period, the sequence exhibits exactly two zeros.
\end{exa}

\section{The Hankel determinants of $G(x,2,k)$ for $k \geq 5$}\label{systemg2r}

\begin{proof}[Proof of the Theorem \ref{thmG2k}]

 Recall that $$G^{2,k}(x,a,b_1,b_2,c)=\frac{a}{1+b_{1} x +b_{2} x^{2}+c x^{k} G^{2,k}(x,a,b_1,b_2,c)},$$ where $k\geq5$.
We apply Proposition \ref{xinu(0,i,ii)} to $G_0:=G^{2,k}(x,a,b_1,b_2,c)$ by repeatedly using the transformation $\tau$.
This results in a shifted periodic continued fractions of order 4 :
\begin{align*}
  G_0(x)& \mathop{\longrightarrow}\limits^\tau G_1^{(1)}(x)\mathop{\longrightarrow}\limits^\tau G^{(1)}_{2}(x)\mathop{\longrightarrow}\limits^\tau G^{(1)}_{3}(x)\mathop{\longrightarrow}\limits^\tau G^{(1)}_{4}(x)\mathop{\longrightarrow}\limits^\tau G^{(1)}_{5}(x)= G_1^{(p+1)}(x)\cdots.
\end{align*}
We obtain
\begin{align}
  H_k(G_0)=a^kH_{k-1}(G_1^{(1)}).\label{g-2-r-G0}
\end{align}

  For $p>1$, computer experiment suggests us to define, for $p \geq 1$.
  $$G_1^{(p)}=\frac{c a x^{k -2}+\left(-p^{2}+p \right) x^{2} b_{2}^{2}+p x b_{1} b_{2}+p b_{2}}{x^{2} G_1^{(p)} +\left(2 p -1\right) b_{2} x^{2}-b_{1} x -1}$$

Apply Proposition \ref{xinu(0,i,ii)} to get $G_2^{(p)}=\tau(G_1^{(p)})$.   This time $d=0$ and $u(x)$ is:
$$u(x)=\frac{\left(2 p -1\right) b_{2} x^{2}-b_{1} x -1}{c a x^{k -2}+\left(-p^{2}+p \right) x^{2} b_{2}^{2}+p x b_{1} b_{2}+p b_{2}} .$$
Since $k \geq 5$, we can decomposition it to $u_L+x^2 u_H$, where
 $u_L(x)=- \frac{1}{pb_2} $ and $$u_H(x)=\frac{p^{2} b_{2}^{2}+x^{k -4} a c}{\left(c a x^{k -2}+(-p^{2} +p)x^{2} b_{2}^{2}+p x b_{1} b_{2}+p b_{2}\right) p b_{2}}.$$
 Then we obtain
$$H_{n}(G_1^{(p)} )= (-p b_2)^n H_{n-1}(G_2^{(p)} ), $$
$$G_2^{(p)} =\frac{c a x^{k -4}}{\left(-c a x^{k -2}+\left(p^{2} b_{2}^{2}-p b_{2}^{2}\right) x^{2}-p x b_{1} b_{2}-p b_{2}\right) x^{2} G_2^{(p)} +2 c a x^{k -2}+p \,x^{2} b_{2}^{2}+p x b_{1} b_{2}+p b_{2}} .$$

Apply Proposition \ref{xinu(0,i,ii)} to get $G_3^{(p)}=\tau(G_2^{(p)})$.   This time $d=k-4$ and $u(x)$ is:
$$ \frac 1{c a}  (2 c a x^{k -2}+p \,x^{2} b_{2}^{2}+p x b_{1} b_{2}+p b_{2}).$$
we can decomposition it to $u_L+x^{k-2} u_H$,
and then we have
 $$H_{n-1}(G_2^{(p)} ) =(-1)^{\binom{k-3}{2}} \left(\frac{c_{0} a_{0}}{p b_{2}}\right)^{n-1} H_{n-k+2}(G_3^{(p)} )$$
 $$ G_3^{(p)}=\frac{a c \left(-c a x^{k -2}+p^{2} x^{2} b_{2}^{2}+p \,x^{2} b_{2}^{2}+p x b_{1} b_{2}+p b_{2}\right)}{p b_{2} \left(x^{k -2} G_3^{(p)}+ p b_{2}+2 c a x^{k -2}-p \,x^{2} b_{2}^{2}-p x b_{1} b_{2}-p b_{2}\right)}.$$

 Apply Proposition \ref{xinu(0,i,ii)} to get $G_4^{(p)}=\tau(G_3^{(p)})$.   This time $d= 0$ and $u(x)$ is:
 $$ -\frac{p b_{2} \left(-2 c a x^{k -2}+p \,x^{2} b_{2}^{2}+p x b_{1} b_{2}+p b_{2}\right)}{a c \left(-c a x^{k -2}+\left(p^{2} b_{2}^{2}+p b_{2}^{2}\right) x^{2}+p x b_{1} b_{2}+p b_{2}\right)} .$$
 Since $k \geq 5$,we can decomposition it to $u_L+x^2 u_H$, where
 $u_L=-\frac{pb_2}{a c}$ and
 $$u_H=\frac{\left(c a x^{k-4}+p^{2} b_{2}^{2}\right) p b_{2}}{\left(-c a x^{k-2}+p^{2} x^{2} b_{2}^{2}+p \,x^{2} b_{2}^{2}+p x b_{1} b_{2}+p b_{2}\right) c a}. $$
 Then we have
 $$H_{n-k+2}(G_3^{(p)})= \left(-\frac{c a}{p b_{2}}\right)^{n-k+2} H_{n-k+1}(G_4^{(p)}) $$
$$G_4^{(p)} = -\frac{p^{2} b_{2}^{2}}{\left(-c a x^{k}+\left(p^{2} b_{2}^{2}+p b_{2}^{2}\right) x^{4}+p \,x^{3} b_{1} b_{2}+b_{2} x^{2} p \right) G_4^{(p)}+\left(-2 p^{2} b_{2}^{2}-p b_{2}^{2}\right) x^{2}-p x b_{1} b_{2}-p b_{2}}.$$
 Apply Proposition \ref{xinu(0,i,ii)} and we have
 $u=\frac{2 b_{2} x^{2} p +b_{2} x^{2}+b_{1} x +1}{p b_{2}}$and
 $u_L =\frac{b_{1} x +1}{p b_{2}}$ and $u_H = \frac{2 b_{2} p +b_{2} }{p b_{2}} $.
 Then we obtain
 $$  H_{n-k+1}(G_4^{(p)})  = (p b_2)^{n-k+1} H_{n-k}  G_1^{(p+1)}.  $$

Combining the above formulas gives the recursion
\begin{align}\label{e-2-r-Gn}
  H_n(G_1^{(p)}) &= (-1)^{\binom{k-2}{2}- k} (c a)^{2n -k +1} H_{n-k}(G_1^{(p+1)})
\end{align}

The initial values are
\begin{align*}
& H_{0}(G^{(n+1)}_{1})=1, \quad H_{1}(G^{(n+1)}_{1})=-(n+1)b_2,\\
&H_{2}(G^{(n+1)}_{1})=H_{3}(G^{(n+1)}_{1})=\cdots=H_{k-3}(G^{(n+1)}_{1})=0,\\
&H_{k-2}(G^{(n+1)}_{1})=(-1)^{\binom{k-3}{2}+k-2}(n+1)b_2(ca)^{k-3}, \\ &H_{k-1}(G^{(n+1)}_{1})=(-1)^{\binom{k-3}{2}+k}(ca)^{k-1}.
\end{align*}
Then the theorem follows by the above initial values, \eqref{g-2-r-G0} and \eqref{e-2-r-Gn}.
\end{proof}

\section{concluding remark}

In this paper, we compute $G^{2,k}(x,a,b_1,b_2,c)$ for $k \geq  4$. For a more general class of lattice paths, the generating function is given by
$$
G^{r,k}(x,a,b_1,b_2,c) = \frac{a}{1 + b_1x + b_2x^2+ \cdots+b_r x^r + cx^k G^{r,k}(x,a,b_1,b_2,c)}.
$$
For the case of $G^{2,4}(x,a,b_1,b_2,c)$, we derive the Hankel determinants through a set of recurrence relations. For the general $G^{r,k}$, where $r$ is a fixed integer and $k \geq 2r+1$, we suggest that the determinant formula computation parallels the case of $G^{2,4}(x,a,b_1,b_2,c)$ for $k \geq 5$ as presented in this paper.

However, when $k \leq 2r$, the problem becomes increasingly challenging and remains unsolved. To our knowledge, for any arbitrary $t$, even if the function $F(x)$ assumes the following simple form,
$$ F(x) = \frac{1}{1 - t x^3 - x^2 F(x)}, $$
the corresponding Hankel determinant remains unknown.\\


\noindent
\textbf{Acknowledgments}
\ \ The authors would like to thank Guoce Xin and Yingrui Zhang for their careful reading and very useful comments.\\

\noindent 
\textbf{Conflict of Interest Statement} 
\ \ The authors declare no conflict of interest.\\

\noindent
\textbf{Data Availability Statement}\ \  Data availability is not applicable to this article as no
new data were created or analyzed in this study.


\end{document}